Two angles are complementary. One angle is 18 degrees less than twice the other. Find the angles.
Answer by checkley77(12844)

x+2x-18=90
3x=90+18
3x=108
x=108/3
x=36 ans. for the smallewr angle.
90-36=54 ans. for the larger angle.
Proof:
36+2*36-18=90
36+72-18=90
90=90